# Probabilistic and Differentiable Wireless Simulation with Geometric Transformers

### v1

@@ Abstract
Modelling the propagation of electromagnetic signals is critical for designing modern communication systems. While there are precise simulators based on ray tracing, they do not lend themselves to solving inverse problems or the integration in an automated design loop. We propose to address these challenges through differentiable neural surrogates that exploit the geometric aspects of the problem. We first introduce the Wireless Geometric Algebra Transformer (Wi-GATr), a generic backbone architecture for simulating wireless propagation in a 3D environment. It uses versatile representations based on geometric algebra and is equivariant with respect to $\mathrm{E}(3)$, the symmetry group of the underlying physics. Second, we study two algorithmic approaches to signal prediction and inverse problems based on differentiable predictive modelling and diffusion models. We show how these let us predict received power, localize receivers, and reconstruct the 3D environment from the received signal. Finally, we introduce two large, geometry-focused datasets of wireless signal propagation in indoor scenes. In experiments, we show that our geometry-forward approach achieves higher-fidelity predictions with less data than various baselines.
@@ 1 Introduction
Modern communication is wireless: more and more, we communicate via electromagnetic waves through the antennas of various devices, leading to progress in and adoption of mobile phones, automotive, AR/VR, and IoT technologies [12, 16]. All these innovations build upon electromagnetic wave propagation. Therefore, modelling and understanding wave propagation in space is a core research area in wireless communication, and remains crucial as we are moving toward new generations of more efficient and spatially-aware wireless technologies.
Wireless signal propagation follows Maxwell’s equations of electromagnetism and is often accurately modelled by state-of-the-art ray-tracing simulation software. However, these simulators take substantial time to evaluate for each scene, cannot be fine-tuned on measurements, and are (usually [29]) not differentiable. This limits their usefulness for solving inverse problems.
In contrast, neural models of signal propagation can be evaluated cheaply, can be trained on real measurements in addition to simulation, and are differentiable and thus well-suited for solving inverse problems. Several such approaches have been proposed recently, often using image-based representations of the inputs and outputs and off-the-shelf vision architectures [35, 51, 44, 6, 23, 46, 52, 34]. However, wireless surrogate modelling faces various challenges. Realistic training data is often scarce, requiring surrogate models to be data efficient. Wireless environments can consist of complex meshes. Finally, input and output data consist of a variety of data types, including the shape of extended 3D objects, point coordinates and spatial orientation of antennas, and information associated with the transmitted signal.
@@
In this work, we present a new approach to modelling wireless signal propagation. It is grounded in the observation that wireless propagation is inherently a geometric problem: a directional signal is transmitted by an oriented transmitting antenna, the signal interacts with surfaces in the environment, and the signal eventually impinges an oriented receiving antenna. We argue that it is critical for neural surrogates to model and flexibly represent geometric aspects (e. g. orientations, shapes) in the propagation environment. We therefore develop surrogate models based on flexible geometric representation and strong geometric inductive biases.
We first propose the Wireless Geometric Algebra Transformer (Wi-GATr), a backbone architecture for wireless signal propagation problems. A key component is a new tokenizer for the diverse, geometric data of wireless scenes. The tokens are processed with a Geometric Algebra Transformer (GATr) network [9]. This architecture is equivariant with respect to the symmetries of wireless channel modelling, but maintains the scalability of a transformer architecture.
Second, we study Wi-GATr models as differentiable, predictive surrogates for the simulator (see Fig. 1a). Here the network predicts observables such as the received power as a function of transmitter position, receiver position, and 3D environment. We show how this enables forward modelling, and in addition, inverse problem solving due to Wi-GATr’s differentiability.
Next, we propose an alternative, more versatile probabilistic approach to prediction and inference tasks: training Wi-GATr diffusion models (Fig. 1b) on the joint distribution of transmitter, receiver, channel information, and 3D environment. At test time, the model can be flexibly conditioned on any available information to predict the received power, localize a transmitter or receiver (Fig. 1c), or even reconstruct the (full or partial) 3D geometry from the wireless signal (Fig. 1d).
To enable machine learning development for wireless problems, we finally introduce two new datasets, Wi3R and WiPTR. Each dataset consists of thousands of indoor scenes of varying complexity and include all the geometric information that characterizes a wireless scene.
Finally, we demonstrate the predictive and the probabilistic models on these datasets. Our experiments show that the Wi-GATr approach gives us a higher-fidelity predictions than various baselines, generalizes robustly to unseen settings, and requires up to 20 times less data for the same performance than a transformer baseline.
@@
2 Background and related work
Wireless signal propagation. How do wireless signals propagate from a transmitting antenna (Tx) to a receiver antenna (Rx) in a (static) 3D environment? While the system is fundamentally described by Maxwell’s equations, for many realistic problems the ray approximation of geometric optics suffices [31]. It approximates the solution to Maxwell’s equations as a sum of planar waves propagating in all directions from Tx. Each planar wave is represented as a ray, characterized by various attributes (e. g., power, phase, delay) since transmission. As a ray reaches an object—that is, it intersects with its mesh—the interaction is modelled as reflection, refraction, or diffraction. During such interactions, the power, phase, polarization, and propagation direction of the wave can change in complex, material-dependent ways. In addition, new rays can emanate from the point of interaction. After multiple interactions, the rays eventually reach the receiving antenna. The Tx and Rx are then linked by a connected path $p$ of multiple rays. The effects on the received signal are described by the channel impulse response (CIR) $h(\tau)=\sum_{p}a_{p}\delta(\tau-\tau_{p})$, where $a_{p}\in\mathbb{C}$ is the complex gain and $\tau_{p}$ the delay of the incoming rays [53].
Maxwell’s equations and in extension ray propagation are highly symmetric. The received signal does not change under rotations, translations, and reflections of the whole scene, as well as the exchange of transmitter and receiver. The latter property is known as reciprocity [37].
@@ 2 Background and related work
Wireless simulators. Wireless propagation models play a key role in design and evaluation of communication systems, for instance by characterizing the gain of competitive designs in realistic settings or by optimizing systems performance as in base station placement for maximal coverage. Statistical approaches [2] represent propagation as a generative model where the parameters of a probabilistic model are fitted to measurements. On the other hand, wireless ray-tracing approaches [1, 5, 29] are increasingly popular due to their high accuracy and because they do not require expensive field measurement collection campaigns.
@@
Neural wireless simulations. Both statistical and ray-tracing simulation techniques are accompanied by their own shortcomings, subsequently mitigated by their neural counterparts. Neural surrogates for statistical models [56, 42, 19, 40] reduce the amount and cost of measurements required. Neural ray tracers [41, 29, 58] address the non-differentiability of simulators using a NeRF-like strategy [38] by parameterizing the scene using a spatial MLP and rendering wireless signals using classic ray-tracing or volumetric techniques. While these techniques are faster than professional ray tracers, they are similarly bottlenecked by expensive bookkeeping and rendering steps (involving thousands of forward passes). In contrast, we propose a framework to simulate wireless signals with a single forward pass through a geometric transformer that is both sample-efficient and generalizes to novel scenes.
Geometric deep learning. The growing field of geometric deep learning [11] aims to incorporate structural properties of a problem into neural network architectures and algorithms. A central concept is equivariance to symmetry groups [15]: a network $f(x)$ is equivariant with respect to a group $G$ if its outputs transform consistently with any symmetry transformation $g\in G$ of the inputs, $f(g\cdot x)=g\cdot f(x)$, where $\cdot$ denotes the group action. Of particular interest to us is the Euclidean group $\mathrm{E}(3)$ of isometries of 3D space, that is, transformations that leave Euclidean distances invariant. This group includes spatial translations, rotations, reflections, and their combinations. As we argued above, the physics of wireless signal propagation are invariant under this group.
@@
GATr. The Geometric Algebra Transformer (GATr) [9] is an $\mathrm{E}(3)$-equivariant architecture for geometric problems. Among equivariant architectures, it stands out in two ways. First, it uses geometric (or Clifford) algebras [22, 14] as representations. For a rigorous introduction to these algebras, we refer the reader to Dorst [20]. From a practical machine learning perspective, these algebras define embeddings for various geometric primitives like 3D points, planes, or $\mathrm{E}(3)$ transformations. We will show that this representation is particularly well-suited for wireless channel modelling. Second, GATr is a transformer architecture [54]. It computes the interactions between multiple tokens through scaled dot-product attention. With efficient backends like FlashAttention [17], the architecture is scalable to large systems, without any restrictions on the sparsity of interactions like in message-passing networks.
Diffusion models. Diffusion models [48, 25, 50] are a class of generative models that iteratively invert a noising process. They have become the de-facto standard in image and video generation [45, 26]. Recently, they have also shown to yield promising results in the generation of spatial and sequential data, such as in planning [30] and puzzle solving [28]. Aside from their generative modelling capabilities, diffusion models provide a flexible way for solving inverse problems [13, 36] through multiplication with an appropriate likelihood term [48]. Furthermore, by combining an invariant prior distribution with an equivariant denoising network, one obtains equivariant diffusion models [33]. These yield a sampling distribution that assigns equal probability to all symmetry transformations of an object, which can improve performance and data efficiency in symmetry problems like molecule generation [27] and planning [10]. We will demonstrate similar benefits in modelling wireless signal propagation.
@@ 3 The Wireless Geometric Algebra Transformer (Wi-GATr)
3.1 Problem formulation
Our goal is to model the interplay between 3D environments, transmitting and receiving antennas, and the resulting transmitted wireless signals. More precisely, we consider wireless scenes consisting of:
  • The 3D geometry $F$ of the environment. We specify it through a triangular mesh with a discrete material class associated with each mesh face.
  • A set of transmitting antennas $t_{i}$ for $i=1,\dots,n_{t}$. Each $t_{i}$ is characterized by a 3D position, an orientation, and any antenna characteristics. We will often focus on the case of a single Tx and then omit the index $i$.
  • Analogously, a set of receiving antennas $r_{i}$ for $i=1,\dots,n_{r}$.
  • The channel or signal $h_{ij}$ between each transmitter $i$ and each receiver $j$, which can be any observable function of the CIR.
In this setting, we consider various downstream tasks:
  • Signal prediction is about predicting the signal received at a single antenna from a single receiver, $p(h|F,t,r)$ with $n_{t}=n_{r}=1$. This is exactly the task that ray-tracing simulators solve. Often, the signal is modelled deterministically as a function $h(F,t,r)$.
  • Receiver localization: inferring the position and properties of a receiving antenna from one or multiple transmitters, $r\sim p(r|F,\{t_{i}\},\{h_{i}\})$, with $n_{r}=1$.
  • Geometry reconstruction or sensing: reconstructing a 3D environment partially, inferring $p(F_{u}|F_{k},t,r,h)$, where $F_{u}$ and $F_{k}$ are the unknown and known subsets of $F$, respectively.
The latter two problems are examples of inverse problems, as they invert the graphical model that simulators are designed for. They are not straightforward to solve with the simulators directly, but we will show how neural surrogates trained on simulator data can solve them.
3.2 Backbone
Core to our approach to this family of inference problems is the Wireless Geometric Algebra Transformer (Wi-GATr) backbone. It consists of a novel tokenizer and a network architecture.
3D environment $F$  Triangular mesh 1 token per mesh face  Mesh face center (point)
 Vertices (points)
 Mesh face plane (oriented plane)
 Material classes  One-hot material emb. (scalars)
Antenna $t_{i}$ / $r_{i}$  Position 1 token per antenna  Position (point)
 Orientation  Orientation (direction)
 Receiving / transmitting  One-hot type embedding (scalars)
 Additional characteristics  Characteristics (scalars)
Channel $h_{ij}$  Antennas 1 token per link  Tx position (point)
 Rx position (point)
 Tx-Rx vector (direction)
 Received power  Normalized power (scalar)
 Phase, delay, …  Additional data (scalars)
@@ 3.2 Backbone
Wireless GA tokenizer. The tokenizer takes as input some subset of the information characterizing a wireless scene and outputs a sequence of tokens that can be processed by the network. A key challenge in the neural modelling of wireless problems is the diversity of types of data involved. As we argued above, a wireless scene consists of the 3D environment mesh $F$, which features three-dimensional objects such as buildings and trees, antennas $t$ and $r$ characterized through a point-like position, an antenna orientation, and additional information about the antenna type, and the characteristics of the channel $h$.
@@
To support all of these data types, we propose a new tokenizer that outputs a sequence of geometric algebra (GA) tokens. Each token consists of a number of elements (channels) of the projective geometric algebra $\mathbb{G}_{3,0,1}$ in addition to the usual unstructured scalar channels. We define the GA precisely in Appendix A. Its main characteristics are that each element is a 16-dimensional vector and can represent various geometric primitives: 3D points including an absolute position, lines, planes, and so on. This richly structured space is ideally suited to represent the different elements encountered in a wireless problem. Our tokenization scheme is specified in Tbl. 1.
@@
Network. After tokenizing, we process the input data with a Geometric Algebra Transformer (GATr) [9]. This architecture naturally operates on our $\mathbb{G}_{3,0,1}$ parameterization of the scene. It is equivariant with respect to permutations of the input tokens as well as $\mathrm{E}(3)$, the symmetry group of translations, rotations, and reflections. These are exactly the symmetries of wireless signal propagation, with one exception: wireless signals have an additional reciprocity symmetry that specifies that the signal is invariant under an role exchange between transmitter and receiver. We will later show how we can incentivize this additional symmetry property through data augmentation. [^1] We also experimented with a reciprocity-equivariant variation of the architecture, but that led to a marginally worse performance without a significant gain in sample efficiency. Finally, because GATr is a transformer, it can process sequences of variable lengths and scales well to systems with many tokens. Both properties are crucial for complex wireless scenes, which can in particular involve a larger number of mesh faces.
3.3 Predictive modelling
The Wi-GATr backbone can be used either in a predictive or probabilistic ansatz. We begin with the predictive modelling of the measured channel information as a function of the complete 3D environment and the information characterizing the transmitter and receiver, $h_{\theta}(F,t,r)$. This regression model is trained in a supervised way on simulated or measured wireless scenes.
Forward prediction. The network thus learns a differentiable, deterministic surrogate for the simulator model $h_{\text{sim}}(F,t,r)$. At test time, we can use the network instead of a simulator to predict the signals in unseen, novel scenes. Compared to a simulator based on ray tracing, it has three advantages: it can be evaluated in microseconds rather than seconds or minutes, it can be finetuned on real measurements, and it is differentiable.

### v2

# Differentiable and Learnable Wireless Simulation with Geometric Transformers

@@ Abstract
Modelling the propagation of electromagnetic wireless signals is critical for designing modern communication systems. Wireless ray tracing simulators model signal propagation based on the 3D geometry and other scene parameters, but their accuracy is fundamentally limited by underlying modelling assumptions and correctness of parameters. In this work, we introduce Wi-GATr, a fully-learnable neural simulation surrogate designed to predict the channel observations based on scene primitives (e. g., surface mesh, antenna position and orientation). Recognizing the inherently geometric nature of these primitives, Wi-GATr leverages an equivariant Geometric Algebra Transformer that operates on a tokenizer specifically tailored for wireless simulation. We evaluate our approach on a range of tasks (i. e., signal strength and delay spread prediction, receiver localization, and geometry reconstruction) and find that Wi-GATr is accurate, fast, sample-efficient, and robust to symmetry-induced transformations. Remarkably, we find our results also translate well to the real world: Wi-GATr demonstrates more than 35% lower error than hybrid techniques, and 70% lower error than a calibrated wireless tracer.
@@ 1 Introduction
Modern communication is wireless: more and more, we communicate via electromagnetic waves through the antennas of various devices, leading to progress in and adoption of mobile phones, automotive, AR/VR, and IoT technologies (Chen et al., 2021; Dahlman et al., 2020). All these innovations build upon electromagnetic (EM) wave propagation. Therefore, modelling and understanding wave propagation in space is a core research area in wireless communication, and remains crucial as we are moving toward new generations of more efficient and spatially-aware wireless technologies.
How can one model the influence of spatial environments on EM waves transmitted by one antenna and received by another? Model-based approaches (specifically wireless ray tracing methods) are popularly used, which identify physical propagation paths and their corresponding characteristics (e. g., power, phase) between the two antennas. The propagation paths are naturally influenced by the environment, causing phenomena such as reflection, diffraction, and transmission. Although popular, model-based have shortcomings: due to their approximate nature, they suffer from a sim-to-real gap in non-trivial environments, and in most cases they are neither learnable nor differentiable.
Hybrid and fully-learnt approaches attempt to address some shortcomings of model-based methods. Hybrid approaches (Orekondy et al., 2022b; Hoydis et al., 2023) retain some aspects of model-based techniques (e. g., evaluating physical propagation paths), and introduce learnable components (e. g., neural evaluation of materials). The learnable parameters help with incorporating real measurement data to reduce the sim-to-real gap. However, similar to model-based approaches, they are once again bottle-necked by availability of accurate and efficient implementations of explicit physical models. In contrast, fully-learnt approaches (Hehn et al., 2023; Levie et al., 2021; Lee & Molisch, 2024), propose an end-to-end neural surrogate that consumes a visual representation of the scene and predicts spatial wireless characteristics. The core idea, which we too leverage in this work, is to enable the neural surrogate to implicitly learn the physical models that best explains the observations. While there have been some advances in this direction, we find the choice of architectures ill-suited and inefficient for the task. Importantly, existing works largely rely on CNN-based architectures and hence are limited to lossy 2D representations of the scene (e.g., top-down binarized image). Furthermore, an image-like representation of the scene prevents them from being differentiable w.r.t relevant simulation parameters (e.g., antenna location and orientation).
Learning neural surrogates directly on 3D data is challenging as the combinations of possible coordinates are infinitely many. Yet, the received wireless signal does not change under rotations, translations, and reflections of the whole scene. To cover these symmetries with data augmentation techniques, it would require sampling from an unbounded space. Instead, by building a model that is $\mathrm{E}(3)$-equivariant one obtains natural generalization to these symmetries.
@@
In this work, we present a new fully-learnt approach towards modelling wireless signal propagation. It is grounded in the observation that wireless propagation is inherently a geometric problem: a directional signal is transmitted by an oriented transmitting antenna, the signal interacts with surfaces in the environment, and the signal eventually impinges an oriented receiving antenna. We argue that it is critical for neural surrogates to model and flexibly represent 3D geometric aspects (e. g., orientations, shapes) in the propagation environment. Therefore, we propose Wi-GATr, a Geometric Algebra Transformer (Brehmer et al., 2023) backbone for surrogate models based on geometric representations and strong geometric inductive biases. A key component is a new tokenizer for the diverse, geometric data of wireless scenes. This architecture is equivariant with respect to the symmetries of wireless propagation, but maintains the scalability of a transformer architecture.
A geometric, fully-learnt treatment of modelling the influence of simulation parameters (e.g., scene mesh, antenna position, orientation) on resulting wireless measurements (e.g., receive signal strength) offers many advantages. First, the underlying inductive biases of Wi-GATr enable sample-efficient learning with the resulting predictions being robust to certain variations. Second, since Wi-GATr is fully differentiable w.r.t. all its simulation parameters, we can leverage it to tackle inverse problems (e.g., localization). Finally, Wi-GATr can be easily adapted into a generative framework and thereby allowing probabilistic inferences.
We present a comprehensive evaluation of Wi-GATr across various tasks and use cases. To facilitate large-scale evaluation and training, we introduce two novel wireless datasets, Wi3R and WiPTR, each comprising thousands of indoor scenes with varying complexity. Our results show that Wi-GATr outperforms competing approaches. For instance, on WiPTR, Wi-GATr achieves an MAE of 0.64 dB using only 10% of the training dataset, surpassing PLViT (1.28 dB) and a Transformer (0.69 dB) trained on the full data. Furthermore, we demonstrate the versatility of our surrogate by leveraging the differentiability of Wi-GATr for receiver localization, achieving accuracy up to 60 cm. Additionally, we show how Wi-GATr can serve as a generative diffusion model trained on joint simulation parameters, enabling the reconstruction of various scene variables, including the intricate mesh geometries. We also train and evaluate Wi-GATr on the real-world measurement dataset DICHASUS (Euchner et al., 2021), where it achieves state-of-the-art results. Specifically, Wi-GATr outperforms both hybrid approaches (with $>$35% reduction in error) and a calibrated wireless ray tracer (with $>$70% reduction in error). Overall, our results indicate that our proposed approach Wi-GATr takes a promising step towards a sample-efficient and robust fully-learnt simulator that can effectively leverage the underlying 3D scene parameters to simulate wireless propagation effects.
@@
2 Related work
@@
Wireless channel models and simulation. Wireless propagation models play a key role in design and evaluation of communication systems, for instance by characterizing the gain of competitive designs in realistic settings or by network performance as in base station placement for maximal coverage. Statistical approaches (3GPP TR 38., 901) represent propagation as a generative model where the parameters of a probabilistic model are fitted to measurements. On the other hand, wireless ray-tracing approaches (Remcom, ; Amiot et al., 2013; Hoydis et al., 2022) approximate wave propagation using geometric optics principles: propagation paths between a transmit and receive antenna are estimated, and corresponding per-path characteristics (e. g., time-of-flight) are calculated. They are sufficiently accurate in many use cases and do not require expensive field measurement collection campaigns.
@@
Neural wireless simulations. Both statistical and ray-tracing simulation techniques are accompanied by their own shortcomings, subsequently mitigated by their neural counterparts. Neural surrogates for statistical models (Ye et al., 2018; O’Shea et al., 2019; Dörner et al., 2020; Orekondy et al., 2022a) reduce the amount and cost of measurements required. Neural ray tracers (Orekondy et al., 2022b; Hoydis et al., 2023; Zhao et al., 2023) address the non-differentiability of simulators using a NeRF-like strategy (Mildenhall et al., 2020) by parameterizing the scene using a spatial MLP and rendering wireless signals using classic ray-tracing or volumetric techniques. While these techniques can be faster than professional ray tracers, they are similarly bottlenecked by expensive bookkeeping and rendering steps (involving thousands of forward passes). In contrast, we propose a framework to simulate wireless signals with a single forward pass through a geometric transformer that is both sample-efficient and generalizes to novel scenes.
@@
Geometric Algebra Transformer. The growing field of geometric deep learning (Bronstein et al., 2021) aims to incorporate structural properties of a problem into neural network architectures and algorithms. A central concept is equivariance to symmetry groups (Cohen, 2021): a network $f(x)$ is equivariant with respect to a group $G$ if its outputs transform consistently with any symmetry transformation $g\in G$ of the inputs, $f(g\cdot x)=g\cdot f(x)$, where $\cdot$ denotes the group action. The Geometric Algebra Transformer (GATr) (Brehmer et al., 2023) is an $\mathrm{E}(3)$-equivariant architecture for geometric problems, where $\mathrm{E}(3)$ is the three dimensional Euclidean group. We will show that this representation is particularly well-suited for wireless channel modelling. Second, GATr is a transformer architecture (Vaswani et al., 2017). It computes the interactions between multiple tokens through scaled dot-product attention. With efficient backends like FlashAttention (Dao et al., 2022), the architecture is scalable to large systems, without any restrictions on the sparsity of interactions like in message-passing networks.
@@ 3 The Wireless Geometric Algebra Transformer (Wi-GATr)
Our goal is to model the wireless channel, i. e., the interplay between the transmitted wireless signal, the 3D environment, and the transmitting and receiving antennas. We first outline the specific problem setup before we provide the details of how we tokenize the environment in geometric algebra primitives that provide the foundation of our Wi-GATr backbone.
3.1 Problem setup
Wireless simulation: fully-learnt forward model. The fundamental problem in simulating wireless propagation is to model an accurate forward process: mapping simulation parameters $\bm{\psi}$ to the ‘wireless channel’ $h$ which describes the transformation between the transmitted signal and the received signal. In this work, we propose a fully-learnable surrogate $h_{\theta}(\bm{\psi})$ for the forward process. The core idea is to implicitly learn physical propagation models that best explain the observations.
Simulation inputs: parameters $\bm{\psi}$. A simulation scenario is defined by a large set of parameters $\bm{\psi}$. Similar to prior works, we consider three groups of representative parameters: (a) scene geometry and materials ${\bm{F}}$: parameterized as a 3D mesh, with each face associated with a discrete material class; (b) transmit antenna properties ${\bm{x}}^{\text{tx}}$: parameterized by 3D position and orientation; (c) receive antenna properties ${\bm{x}}^{\text{rx}}$: also parameterized by 3D position and orientation. Note that in our setting, we assume that antenna radiation patterns are shared between all receivers and transmitters, respectively, and as a consequence, they need to be learned implicitly. To the best of our knowledge, we present the first fully-learnt surrogate that works on the full 3D geometry (as opposed to 2D/2.5D image representations).
Simulation outputs: wireless channel $h$. For a fixed set of simulation inputs, we want the simulator to predict the wireless channel $h$. Since we propose a learnable model, we rely on channel measurements for supervised training which can be collected using conventional devices (e. g., mobile phones). Measuring the exact channel $h$ is challenging (Euchner et al., 2021), which is why we describe the signal transformation by averaged statistics of the channel. Similar to prior works, we model real-valued, scalar channel characteristics $h$: non-coherent received power, band-limited received power, and band-limited delay spread.
3.2 Wi-GATr backbone
We want a learnable forward-model for simulations $h_{\theta}(\bm{\psi})$, that can effectively leverage a wireless scene $\bm{\psi}$ to reason about spatially-varying channels $h$ (e. g., receive power). Our core insight is that the forward-process is inherently geometric: electromagnetic waves are emitted by an oriented transmit antenna, interact (e. g., reflect and refract) multiple times with various surfaces in space and finally reach an oriented receive antenna at a different location. To exploit this inherent geometric nature of propagation, we propose ‘Wi-GATr’, a Wireless Geometric Algebra Transformer (Wi-GATr) backbone. Wi-GATr consists of two main components: a tokenizer and a network architecture. The tokenizer embeds the information of the wireless scene into geometric algebra while the network learns to model the channel. We now elaborate on these individual components.
@@ 3.2 Wi-GATr backbone
Wireless GA tokenizer. The tokenizer takes as input information characterizing simulation parameters $\bm{\psi}$ and outputs a sequence of tokens that can be processed by the network. A key challenge of the tokenizer is efficiently representing a variable set of diverse data types. In our case, the data types range from 3D environment mesh ${\bm{F}}$, which features three-dimensional objects with dielectric material properties such as buildings, to antennas ${\bm{x}}^{\text{tx}}$ and ${\bm{x}}^{\text{rx}}$ characterized through a point-like position, an antenna orientation, and additional information about the antenna type, and the characteristics of the channel $h$.
@@
To support all of these data types, we propose a new tokenizer that outputs a sequence of geometric algebra (GA) tokens. Each token consists of a number of elements (channels) of the projective geometric algebra $\mathbb{G}_{3,0,1}$ in addition to the usual unstructured scalar channels. We define the GA precisely in Appendix A. Its main characteristics are that each element is a 16-dimensional vector and can represent various geometric primitives: 3D points including an absolute position, lines, planes, and so on. The richly structured space of projective geometric algebra $\mathbb{G}_{3,0,1}$ is thus ideally suited to represent the different elements encountered in a wireless problem. It naturally supports scalars for wireless scalar fields (e. g., power, delay spread), points for positions (e. g., antennas, mesh vertices), vectors for orientations (e. g., antennas), and planes for surfaces (e. g., mesh faces). Details of our tokenization scheme are specified in Appenndix B.
@@
Network. After tokenizing, we process the input data with a Geometric Algebra Transformer (GATr) (Brehmer et al., 2023). This architecture naturally operates on our $\mathbb{G}_{3,0,1}$ parameterization of the scene. It is equivariant with respect to permutations of the input tokens as well as $\mathrm{E}(3)$, the symmetry group of translations, rotations, and reflections. These are exactly the symmetries of wireless signal propagation, with one addition: wireless signals have an additional reciprocity symmetry that specifies that the signal is invariant under an role exchange between transmitter and receiver. We will later show how we can incentivize this additional symmetry property through data augmentation. [^1] We also experimented with a reciprocity-equivariant variation of the architecture, but that led to a marginally worse performance without a significant gain in sample efficiency. Finally, because GATr is a transformer, it can process sequences of variable lengths and scales well to systems with many tokens. Both properties are crucial for complex wireless scenes, which can in particular involve a larger number of mesh faces.
3.3 Applications beyond forward modelling
In the previous section, we proposed a Wi-GATr backbone for forward-simulations $h_{\theta}(\bm{\psi})$. We now highlight two additional benefits of the proposed backbone. First, exploiting differentiability of $h_{\theta}$ to solve inverse problems. Second, adapting the Wi-GATr within a generative model formulation and thereby allowing probabilistic inferences (either for forward, or inverse problems).